-6=-(x-5)-3(5+2x)-4(2x-4)
Answer by rfer(10417)

-6=-(x-5)-3(5+2x)-4(2x-4)
-6=-x+5-15-6x-8x+16
-6=6-15x
-12=-15x
x=12/15
x=4/5